# Forecasting-Based Biomedical Time-series Data Synthesis for Open Data and Robust AI

### v1

@@ 3.1 Data Preparation
The EEG dataset underwent a two-stage labeling process to achieve accurate sleep stage identification. Initially, EMG signals served as reference signals to differentiate between SLEEP and WAKE states by comparing EMG power against a baseline threshold. Subsequently, epochs identified as SLEEP underwent EEG signal analysis using frequency domain filtering via fast Fourier transform (FFT) to characterize sleep stages with high specificity. We intentionally filtered frequency bands to highlight pronounced peaks within the delta frequency range (0.5-4 Hz) for NREM sleep and distinct peaks within the theta frequency range (4-8 Hz) for REM sleep. This labeled EEG dataset provided the foundation for training and evaluating synthetic data generation using time-series forecasting models.
@@ 3.2 Performance Analysis
We evaluated 16 state-of-the-art time-series forecasting models for synthetic biomedical data generation across four subjects (Table LABEL:tab:1). Models in the Transformer family (SOFTS, TiDE, iTransformer, TFT, PatchTST, TimesNet) consistently achieved the highest accuracy across all subjects, demonstrating strong capacity to model long-range dependencies inherent in biomedical signals. MLP-based models (N-BEATS, NBEATSx, N-HiTS, DLinear, DeepNPTS) also exhibited robust performance, particularly at larger window sizes, highlighting their ability to capture essential signal structures. RNN-based models (DilatedRNN, DeepAR) contributed positively, though their performance varied with subject characteristics. CNN-based models (TCN, BiTCN) showed competitive accuracy in several settings. Additionally, KAN, belonging to the ‘any’ category, ranked among the top performers in multiple cases. The synthetic-only (S) training condition frequently approached original-only accuracy, indicating the high fidelity of the generated data. Thus, the open release of these synthetic datasets through platforms such as Hugging Face and Kaggle is expected to further accelerate progress in biomedical AI.
@@
Additionally, the accuracy patterns across varying window sizes and model families reveal further insights (Fig. LABEL:fig:3). In the S condition, larger window sizes generally yielded higher accuracy, with Transformer and MLP models maintaining strong and consistent performance across settings. The CNN and RNN models exhibited greater sensitivity to window size, reflecting their dependence on local temporal patterns. Under the O+S condition, accuracy improved substantially across all subjects, and variability across window sizes was significantly reduced.
@@ 3.3 Class-wise Performance Analysis
3.4 Visualization of Original vs. Synthetic Data.
@@
We perform UMAP-based visualization [36] to compare the distribution of original and synthetic data across all subjects (Fig. 5). For each subject, 100 samples per class were projected into a 2D space, revealing distinct clusters for WAKE, NREM, and REM stages. In the original data (O), clusters exhibit clear separation across all subjects (Fig. 5a), indicating well-preserved inter-class distinctions. Furthermore, local neighborhood structures remained consistent, reflecting the intrinsic temporal characteristics of EEG signals. The slight overlaps in Subject 3 between NREM and REM were observed, likely due to class imbalance effects. These original UMAP projections provide a robust reference for evaluating synthetic data fidelity and preserving class boundaries in biomedical time-series synthesis.
@@
We then visualized the synthetic-only (S) data using the best-performing generator for each subject (Fig. 5b). Transformer-based models such as SOFTS and TiDE, and forecasters such as KAN produced synthetic clusters that closely aligned with their original counterparts. In Subjects 1 and 3, WAKE and NREM clusters showed minimal shift, suggesting that key temporal patterns were effectively captured. Synthetic REM clusters remained somewhat more dispersed across subjects, introducing diversity without compromising core class identity. Subject 4’s synthetic NREM points displayed broader spread, reflecting the generator’s effort to enrich under-represented regions. No spurious clusters or artifacts were observed, confirming the physiological plausibility of the synthetic data. The results indicate that time-series forecasters can produce synthetic EEG data that respects both global structure and local variations of the original signals.
@@
Finally, we examined the combined UMAP plots with original and synthetic data merged (O+S) (Fig. 5c). In all subjects, synthetic data effectively complemented the original manifold, filling sparse regions and enhancing data coverage. Notably, in Subject 3, synthetic REM points extended into low-density areas, addressing feature space gaps. Moreover, combined clusters exhibited tighter boundaries, particularly for NREM in Subject 4, indicating improved class balance and coverage. Transformer-based and MLP-based generators such as DeepAR, DLinear, and TiDE contributed prominently to this harmonization across latent space. Thus, the insights underscore the utility of synthetic data in enhancing diversity and generalization for downstream biomedical AI tasks.
@@ 3.5 Comparison with GAN-based Synthesis
We compare the performance of synthetic data generated by TimeGAN [37], a representative generative model for time-series data, with that of our proposed time-series forecaster-based approach (Fig. 6). Under the synthetic-only (S) condition, our forecaster-based method consistently outperformed TimeGAN across all subjects. For instance, in Subject 1, TimeGAN achieved 42.00%, whereas our method reached 91.00% with SOFTS. Similarly, in Subject 2, TimeGAN yielded 49.47%, while our TiDE-based approach achieved 96.76%. In Subjects 3 and 4, the gap remained substantial, with our models exceeding 95% and 82%, respectively, compared to TimeGAN’s 39.45% and 32.96%. These results indicate that our forecasting models more effectively capture temporal dynamics critical for biomedical time-series data. Moreover, the low variance observed in our method highlights its robustness across different seeds.
@@ 4 Discussion
This study shows that forecasting-based synthetic data generation is highly effective for biomedical time-series applications. Across subjects and conditions, our approach consistently surpassed the widely used TimeGAN method, and synthetic data from state-of-the-art forecasters preserved class-specific characteristics while enhancing model generalization when combined with original data. Moreover, UMAP visualizations confirmed close alignment with original distributions while enriching sparse regions. In addition, the method remained stable across seeds and window sizes, underscoring its robustness. Thus, forecasting-based synthesis emerges as a scalable and practical solution for privacy-preserving biomedical data augmentation.
Although this work focused on dataset generation, the optimal quantity of synthetic data for diverse applications remains unresolved, and clarifying how different synthetic-to-original ratios influence performance is a key challenge. Furthermore, determining when additional augmentation yields diminishing returns is essential for efficient dataset design. While balanced O+S combinations produced consistent gains here, task-specific requirements may call for different ratios. Therefore, adaptive strategies that adjust generation based on model feedback could further improve outcomes. Looking ahead, future efforts should pursue these directions to establish efficient, application-specific synthesis protocols that maximize utility while controlling computational and privacy costs.
@@ 5 Conclusion
In this work, we demonstrate that forecasting-based synthetic data generation offers a powerful and scalable approach for biomedical time-series applications. Our method consistently outperformed conventional GAN-based synthesis, producing high-fidelity data that enhances AI model robustness. Moreover, the open-source release of these synthetic datasets will help democratize access to biomedical data while preserving patient privacy. Importantly, our results show that combining synthetic with original data improves generalization and enables more reliable clinical AI systems. Looking ahead, optimizing the quantity and composition of synthetic data will further maximize its impact on open-source contributions and robust biomedical AI development.
@@ Acknowledgments
This research was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP)-ITRC (Information Technology Research Center) grant funded by the Korea government (MSIT) (IITP-2025-RS-2020-II201787).

### v2

@@ 3.1 Data Preparation
The EEG dataset underwent a two-stage labeling process to achieve accurate sleep stage identification following methods described in [jo2022general] (Fig. 3). Initially, EMG signals served as reference signals to differentiate between SLEEP and WAKE states by comparing EMG power against a baseline threshold. Subsequently, epochs identified as SLEEP underwent EEG signal analysis using frequency domain filtering via fast Fourier transform (FFT) to characterize sleep stages with high specificity. We intentionally filtered frequency bands to highlight pronounced peaks within the delta frequency range (0.5-4 Hz) for NREM sleep and distinct peaks within the theta frequency range (4-8 Hz) for REM sleep.
In particular, NREM sleep is characterized by an increase in delta wave activity, while strong theta wave power is observed during REM sleep. Then, we filtered the frequency bands to highlight pronounced peaks within the delta frequency range (0.5-4 Hz) for NREM sleep and distinct peaks within the theta frequency range (4-8 Hz) for REM sleep. This labeled EEG dataset provide the foundation for synthetic data generation using time-series forecasting models.
@@ 3.2 Performance Analysis
We evaluated 16 state-of-the-art time-series forecasting models for synthetic biomedical data generation across four subjects (Table LABEL:tab:1). The models in the Transformer family (SOFTS, TiDE, iTransformer, TFT, PatchTST, TimesNet) consistently achieved the highest accuracy across all subjects, demonstrating strong capacity to model long-range dependencies inherent in biomedical signals. As [trirat2024universal], this suggests that the global receptive field of self-attention mechanisms is well suited for modeling long-range dependencies. Notably, MLP-based models (N-BEATS, NBEATSx, N-HiTS, DLinear, DeepNPTS) also exhibited robust performance, particularly at larger window sizes, highlighting their ability to capture essential signal structures. Meanwhile, RNN-based models (DilatedRNN, DeepAR) contributed positively, though their performance varied with subject characteristics. The CNN-based models (TCN, BiTCN) showed competitive accuracy in several settings. Additionally, KAN, belonging to the ‘any’ category, ranked among the top performers in multiple cases. The synthetic-only (S) training condition frequently approached original-only accuracy, indicating the high fidelity of the generated data. Thus, the open release of these synthetic datasets through platforms such as Hugging Face and Kaggle is expected to further accelerate progress in biomedical AI.
@@
In addition, the test accuracy patterns across varying window sizes and model families reveal further insights (Fig. LABEL:fig:3). In the S condition, larger window sizes generally yielded higher accuracy, with Transformer and MLP models maintaining strong and consistent performance across settings. By contrast, the CNN and RNN models exhibited greater sensitivity to window size, reflecting their dependence on local temporal patterns. Under the O+S condition, the test accuracy improved substantially across all subjects, and variability across window sizes was significantly reduced.
@@ 3.3 Class-wise Performance Analysis
3.4 Visualization of Original vs. Synthetic Data
@@
We perform UMAP-based visualization [McInnes2018] to compare the distribution of original and synthetic data across all subjects (Fig. 5). The UMAP preserves manifold topology, allowing spatial alignment and inter-class separation to reliably indicate feature fidelity and decision boundary retention, respectively [ghojogh2023elements]. Note that the synthetic data were generated using each subject’s best-performing forecaster (S / O+S): SOFTS / DeepAR for Subject 1, TiDE / TiDE for Subject 2, KAN / TiDE for Subject 3, and SOFTS / DLinear for Subject 4, which are consistently applied in all subsequent experiments. For each subject, 100 samples per class were projected into a 2D space, revealing distinct clusters for WAKE, NREM, and REM stages. In the original data (O), clusters exhibit clear separation across all subjects (Fig. 5a), indicating well-preserved inter-class distinctions except Subject 2. Note that slight overlaps between NREM and WAKE were observed in Subject 2, reflecting the inherent low separability between these classes specific to this subject. Moreover, local neighborhood structures remained consistent, reflecting the intrinsic temporal characteristics of EEG signals. These original UMAP projections provide a robust reference for evaluating synthetic data fidelity and preserving class boundaries in biomedical time-series synthesis.
@@
We then visualized the synthetic-only (S) data using the best-performing generator for each subject (Fig. 5b). The Transformer-based models such as SOFTS and TiDE, and forecasters such as KAN, produced synthetic clusters that closely aligned with their original counterparts across all subjects. In Subjects 1–3, the WAKE and NREM clusters remained compact with minimal drift, indicating that key temporal dynamics were effectively captured. Notably, Subject 4 demonstrated even tighter clustering across all classes, suggesting that the generator concentrated more heavily on dominant and reliable feature regions. No spurious clusters or artifacts were observed, confirming the physiological plausibility of the synthetic data. The results indicate that time-series forecasters can produce synthetic EEG data that respects both global structure and local variations of the original signals.
@@
Finally, we examine UMAP plots with original and synthetic data merged (O+S) (Fig. 5c). In all subjects, synthetic data effectively complemented the original manifold, filling sparse regions and enhancing data coverage. Notably, in Subject 1, synthetic REM points extended into low-density areas, addressing feature space gaps. Moreover, except for Subject 1, the combined clusters exhibit tighter boundaries, indicating improved class balance and coverage. Thus, the insights underscore the utility of synthetic data in enhancing diversity and generalization for downstream biomedical AI tasks.
3.5 Evaluation of Quality and Privacy Metrics
Table 1: Quantitative assessment of synthetic data quality and privacy preservation across subjects. The table presents quality metrics and privacy metrics for Synthetic-only (S) and Combined (O+S) modes.
Subject Mode Quality Metrics Privacy Metrics
Stat Sim. Corr Sim. DCR NNDR
Subject 1 S 0.9667 0.9639 0.6592 0.9569
O+S 0.9731 0.9567 0.4558 0.9759
Subject 2 S 0.9744 0.9504 0.3619 0.9525
O+S 0.9735 0.9460 0.3570 0.9598
Subject 3 S 0.9742 0.9747 0.4571 0.9676
O+S 0.9750 0.9866 0.4430 0.9766
Subject 4 S 0.9711 0.9763 0.4951 0.9266
O+S 0.9779 0.9797 0.1804 0.8942
To evaluate both the quality and privacy of the generated datasets, we conducted a quantitative assessment using the best-performing forecaster for each subject under the synthetic-only (S) and combined (O+S) settings (Table 1). We randomly sampled 50 synthetic instances per condition to compute the quality metrics—Statistical Similarity (Stat Sim.) and Correlation Similarity (Corr Sim.) [yang2024structured]. Using the same samples, we evaluated the privacy metrics, Distance to Closest Record (DCR) and Nearest Neighbor Distance Ratio (NNDR) [steier2025synthetic]. The quality metrics show that the synthetic data consistently preserved key temporal and statistical characteristics, with Stat Sim. values exceeding 0.96 across all subjects.
In terms of privacy preservation, we analyzed distance-based metrics to assess the risk of the synthetic samples simply memorizing the training data. The DCR values were non-zero across all subjects in both the S and O+S settings, indicating measurable separation from the real data and thereby reducing privacy risk. In addition, the NNDR values remained consistently high across subjects, with Subject 3 reaching 0.9676 in the S mode, indicating strong resistance to nearest-neighbor–based re-identification. Overall, the results show that the method yields representative data without duplicating the originals. Note that this evaluation uses a vanilla setting, and adding differential privacy [abadi2016deep] can further strengthen privacy.
@@ 3.6 Comparison with GAN-based Synthesis
We compare the performance of synthetic data generated by TimeGAN [yoon2019time], a representative generative model for time-series data, with that of our proposed time-series forecaster-based approach (Fig. 6). Under the synthetic-only (S) condition, our forecaster-based method consistently outperformed TimeGAN across all subjects. For instance, in Subject 1, TimeGAN achieved 42.00%, whereas our method reached 91.00% with SOFTS. Similarly, in Subject 2, TimeGAN yielded 49.47%, while our TiDE-based approach achieved 96.76%. In Subjects 3 and 4, the gap remained substantial, with our models exceeding 95% and 82%, respectively, compared to TimeGAN’s 39.45% and 32.96%. These results indicate that our forecasting models more effectively capture temporal dynamics critical for biomedical time-series data. Moreover, the low variance observed in our method highlights its robustness across different seeds.
@@ 4 Discussion
Figure 7: Impact of synthetic data ratio on classification accuracy. Test accuracy for Synthetic-only (S, green triangles) and Combined (O+S, blue squares) conditions is plotted against synthetic proportions (0.1–1.0), relative to the Original-only baseline (O, black dashed line). Shaded regions represent the standard deviation across five independent seeds.
To evaluate the impact of synthetic data volume, we assessed model performance across synthetic-to-original ratios ranging from 0.1 to 1.0 (Fig. 7). In the synthetic-only (S) condition, accuracy was highly sensitive to data volume. For example, Subject 4 improved from 53.46% to 82.08% as the ratio increased, indicating that sufficiently large synthetic datasets are required to capture the underlying distribution. By contrast, the combined (O+S) condition showed stable robustness and consistent gains over the baseline, yet accuracy tended to saturate beyond a ratio of 0.8. This suggests that adopting adaptive generation strategies that adjust synthetic volume based on model feedback may yield additional improvements. Note that although this work demonstrates the utility of the proposed framework for EEG signals, further validation is needed to extend it to more complex modalities such as multi-channel or irregular biomedical time-series. Looking ahead, future work should explore these directions to establish application-specific synthesis protocols that maximize utility while managing computational and privacy costs.
@@ 5 Conclusion
In this work, we demonstrate that forecasting-based synthetic data generation offers a powerful and scalable approach for biomedical time-series applications. Our method consistently outperformed conventional GAN-based synthesis, producing high-fidelity data that enhances AI model robustness. Moreover, the open-source release of these synthetic datasets will help democratize access to biomedical data while preserving patient privacy. Importantly, our results show that combining synthetic with original data improves generalization and enables more reliable clinical AI systems. In future work, we aim to develop this framework for specific clinical pathologies and integrate synthetic pipelines into real-world healthcare workflows.
@@ Acknowledgments
This research was partly supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP)-ITRC (Information Technology Research Center) grant funded by the Korea government (MSIT) (IITP-2026-RS-2020-II201787, contribution rate: 50%). Also, this research was supported in part by a grant of the Korea-US Collaborative Research Fund (KUCRF), funded by the Ministry of Science and ICT and Ministry of Health & Welfare, Republic of Korea (grant number: RS-2025-16022980, contribution rate: 50%).